OpenStudy (anonymous):

Simplify the complex fraction. x divide y x=4/9, 5 1/5

OpenStudy (anonymous):

Is y= 5 1/5?

OpenStudy (anonymous):

yes

OpenStudy (dumbcow):

\[\frac{\frac{4}{9}}{\frac{26}{5}} = \frac{4}{9}*\frac{5}{26} = \frac{10}{117}\]

OpenStudy (anonymous):

\[\frac{\left(\frac{4}{9}\right)}{\left(\frac{26}{5}\right)}=\frac{10}{117} \]

OpenStudy (anonymous):

And robtobey finishes 2nd. Congratulations dumbcow.

OpenStudy (dumbcow):

shoot, are we racing?? haha dang i need to learn to type faster then for future

OpenStudy (anonymous):

I can't type and keep my eye on the current postings.

OpenStudy (anonymous):

thanks guys

OpenStudy (dumbcow):

your welcome